A delivery truck travels 21 m north, 16 m east, and 26 m south. What is it's final displacement from the origin?
1
Expert's answer
2020-10-19T13:24:27-0400
"\vec{s}=(0,-21)+(16,0)+(0,26)=(16,5)\ m"

"s=\sqrt{16^2+5^2}=16.8\ m"

"\theta=\arctan{\frac{5}{16}}=17.35\degree\text{S of E}"